At what temperature would the average speed of helium atoms equal the escape speed from the Earth, 1.12 x 10^4 m/s, the mass of a helium atom being = 6.64310227 kg
1
Expert's answer
2021-02-07T19:17:26-0500

The average speed of helium atoms can be found from the formula:


"v_{avg,He}=\sqrt{\dfrac{8k_BT}{\pi m}}."

In order to find the temperature at which the average speed of helium atoms equals the escape speed from the Earth we must equate "v_{avg,He}" and "v_{esc}":


"v_{esc}=\sqrt{\dfrac{8k_BT}{\pi m}},""T=\dfrac{\pi mv_{esc}^2}{8k_B},""T=\dfrac{\pi\cdot6.64\cdot10^{-27}\ kg\cdot(1.12\cdot10^4\ \dfrac{m}{s})^2}{8\cdot1.38\cdot10^{-23}\ \dfrac{J}{K}}=2.37\cdot10^4\ K."